# Fractional Infinity Laplacian with Obstacle

### v1

@@ 3.2. Perron’s Method
$-L[g(x_{0})-C\,\psi_{\beta,x_{0}}]+f(g(x_{0})-C\,\psi_{\beta,x_{0}})\leq C\,\text{{\rm diam}}(\Omega)^{\beta-\alpha}\left(\dfrac{r_{\star}^{\beta}-1}{(r_{\star}-1)^{\alpha}}-1\right)+f(||g||_{\infty})\leq 0$
as soon as
$C\geq\frac{\mbox{diam}(\Omega)^{\alpha-\beta}\,f(||g||_{\infty})}{\bigg{[}1-\dfrac{r_{\star}^{\beta}-1}{(r_{\star}-1)^{\alpha}}\bigg{]}}.$
Moreover, it is clear that $u^{-}\in C^{0,\beta}({\Omega})$. Thanks to Proposition 3.6, this yields that $u^{-}$ is a subsolution. On the other hand, it is clear that
$u^{-}(x_{0})\geq g(x_{0})-C\,\psi_{\beta,x_{0}}(x_{0})=g(x_{0}),\qquad\mbox{for all}\,\,\,x_{0}\in\partial\Omega.$
Thanks to the $\beta-$Hölder regularity of $g$ on $\partial\Omega$, we have the following estimate for $C\geq[g]_{\alpha,\partial\Omega}$
$g(x_{0})-C\,\psi_{\beta,x_{0}}(x)\leq g(x),\qquad\mbox{for all}\,\,\,x_{0}\in\partial\Omega.$
Hence,
$u^{-}(x)\leq g(x),\qquad\mbox{for all}\,\,\,x\in\partial\Omega.$
Consequently, $u^{-}=g$  on  $\partial\Omega$.
Now, we define
$u^{+}(x)=\min\{g(x_{0})+C\,\psi_{\beta,x_{0}}(x),\,x_{0}\in\partial\Omega\}\qquad\mbox{for all}\,\,\,x\in\overline{\Omega}.$
One has
$-L[g(x_{0})+C\,\psi_{\beta,x_{0}}]+f(g(x_{0})+C\,\psi_{\beta,x_{0}})\geq-C\,L[\psi_{\beta,x_{0}}]+f(-||g||_{\infty})\geq 0$
provided that $C$ is large enough. Hence, thanks to Remark 3.7, $u^{+}\in C^{0,\beta}({\Omega})$ is a viscosity supersolution. In addition, $u^{+}=g$  on  $\partial\Omega$.
Finally, let us check that $u^{-}\leq u^{+}$  on  $\overline{\Omega}$. For  $x_{0},\,x_{1}\in\partial\Omega$, we have from Lemma 2.1 that
$g(x_{0})-g(x_{1})\leq C|x_{1}-x_{0}|^{\beta}\leq C(|x-x_{0}|^{\beta}+|x-x_{1}|^{\beta}),\qquad\mbox{for every}\,\,\,x\in\overline{\Omega}.$
Fix $x\in\overline{\Omega}$. Hence, one has
$g(x_{0})-C\,\psi_{\beta,x_{0}}(x)\leq g(x_{1})+C\,\psi_{\beta,x_{1}}(x),\qquad\mbox{for every}\,\,\,\,x_{1}\in\partial\Omega.$
Thus,
$g(x_{0})-C\,\psi_{\beta,x_{0}}(x)\leq\min\{g(x_{1})+C\,\psi_{\beta,x_{1}}(x),\,\,x_{1}\in\partial\Omega\}=u^{+}(x),\qquad\mbox{for every}\,\,\,x_{0}\in\partial\Omega.$
Consequently, we get that
$u^{-}(x)=\sup\{g(x_{0})-C\,\psi_{\beta,x_{0}}(x),\,\,x_{0}\in\partial\Omega\}\leq u^{+}(x)\qquad\mbox{on}\,\,\,\overline{\Omega}.$
Yet, this concludes the proof.
We are now ready to prove Theorem 1.1. We first show the regularity of subsolutions and then prove the existence result.
@@
Assume $f$ is non-decreasing. If $u$ is a bounded continuous viscosity subsolution of (1.0.8), then $u$ is locally $\beta-$Hölderian, for any $\beta<\alpha$. More precisely, we have
$[u]_{C^{0,\beta}(\omega)}\leq\max\bigg{\{}\frac{2||u||_{\infty}}{\text{dist}(\omega,\partial\Omega)^{\beta}},\frac{[\mbox{diam}(\Omega)]^{\alpha-\beta}\,[f(-||u||_{\infty})]_{-}}{1-\Psi(r_{\star})}\bigg{\}},\qquad\mbox{for every}\,\,\,w\subset\subset\Omega,$
@@
In addition, assume that $g\in C^{0,\beta}(\partial\Omega)$, if $u$ is a continuous viscosity solution of (1.0.8) with $u=g$ on $\partial\Omega$, then $u$ is $\beta-$Hölderian in $\overline{\Omega}$, and we have the following estimate:
$||u||_{C^{0,\beta}(\overline{\Omega})}\leq C\bigg{(}\alpha,\,\beta,\,\mbox{diam}(\Omega),\,||g||_{C^{0,\beta}(\partial\Omega)},\,[f(\pm||g||_{\infty})]_{\pm},\,[f(-||u||_{\infty})]_{-}\bigg{)}.$
@@
Fix $x_{0}\in\omega\subset\subset\Omega$. Thanks to Proposition 3.4, $u$ is a viscosity subsolution of (1.0.8) on $\Omega\backslash\{x_{0}\}$. Now, we define
@@
Recalling Proposition 2.2, $v$ is a strict supersolution on $\Omega\backslash\{x_{0}\}$ since
$-L[u(x_{0})+C|x-x_{0}|^{\beta}]+f(u(x_{0})+C|x-x_{0}|^{\beta})\geq-CL[|x-x_{0}|^{\beta}]+f(-||u||_{\infty})>0$
as soon as $C\geq\frac{[\mbox{diam}(\Omega)]^{\alpha-\beta}\,[f(-||u||_{\infty})]_{-}}{1-\Psi(r_{\star})}$. Moreover, one has $v(x_{0})=u(x_{0})$, and for every $x\in\partial\Omega$, we have
$u(x)-v(x)=u(x)-u(x_{0})-C|x-x_{0}|^{\beta}\leq 2||u||_{\infty}-C\,\text{dist}(\omega,\partial\Omega)^{\beta}\leq 0$
as soon as we choose the constant $C\geq 2||u||_{\infty}/\text{dist}(\omega,\partial\Omega)^{\beta}$. Thanks to the comparison principle in Proposition 3.5, we infer that $u<v$ in $\Omega\backslash\{x_{0}\}$ and so, we have
@@
$-L[w^{+}]+f(w^{+})$ $=-C\,L[\psi_{\beta,x_{0}}]+f(g(x_{0})+C\,|x-x_{0}|^{\beta})\geq C|x-x_{0}|^{\beta-\alpha}[1-\Psi(r_{\star})]+f(-||g||_{\infty})>0$
@@
$C\geq\frac{\text{{\rm diam}}(\Omega)^{\alpha-\beta}\,[f(-||g||_{\infty})]_{-}}{1-\Psi(r_{\star})}.$
@@
$-L[w^{-}]+f(w^{-})=C\,L[\psi_{\beta,x_{0}}]+f(g(x_{0})-C\,|x-x_{0}|^{\beta})\leq C|x-x_{0}|^{\beta-\alpha}[\Psi(r_{\star})-1]+f(||g||_{\infty})<0$
@@
$C\geq\frac{\text{{\rm diam}}(\Omega)^{\alpha-\beta}\,[f(||g||_{\infty})]_{+}}{1-\Psi(r_{\star})}.$
@@
Under the assumptions that $f$ is non-decreasing, continuous and $g\in C^{0,\beta}(\partial\Omega)$ for some $\beta>0$, there exists a viscosity solution $u$ to Problem (1.0.8).
@@
Without loss of generality, we assume that $\beta<\alpha$. Set
@@
$||w||_{\infty}\leq\max\{||u^{+}||_{\infty},||u^{-}||_{\infty}\}:=\Lambda.$
Then, by Proposition 3.9, one has
$[w]_{C^{0,\beta}(\omega)}\leq\max\bigg{\{}\frac{2\Lambda}{\text{dist}(\omega,\partial\Omega)^{\beta}},\frac{[\mbox{diam}(\Omega)]^{\alpha-\beta}\,[f(-\Lambda)]_{-}}{1-\Psi(r_{\star})}\bigg{\}},\qquad\mbox{for every}\,\,\,w\subset\subset\Omega.$
Hence, we infer that $u$ is locally $\beta-$Hölder. In particular, $u$ is continuous in $\Omega$. From Proposition 3.6, this implies that $u$ is a subsolution of (1.0.8) with $u=g$ on $\partial\Omega$.
@@
$-L[\varphi_{\delta}](x_{0})+f(\varphi_{\delta}(x_{0}))\leq-L[\varphi](x_{0})+C\delta+f(\varphi(x_{0}))<0,$
@@
$-L[\varphi_{\zeta}](x_{\zeta})+f(\varphi_{\zeta}(x_{\zeta}))>0.$ (3.2.2)
@@
$-L[\varphi_{\zeta}](x_{\zeta})+f(\varphi_{\zeta}(x_{\zeta}))\leq 0,$
@@
$-L[\varphi_{\zeta}](x_{\zeta})+f(\varphi(x_{\zeta}))\leq-L[\varphi](x_{\zeta})+f(\varphi(x_{\zeta})).$
@@
$[-L[\varphi_{\zeta}](x_{\zeta})+f(\varphi_{\zeta}(x_{\zeta}))]+[f(\varphi(x_{\zeta}))-f(\varphi_{\zeta}(x_{\zeta}))]\leq-L[\varphi](x_{\zeta})+f(\varphi(x_{\zeta})).$ (3.2.3)
@@
Assume the boundary datum $g\geq 0$ on $\partial\Omega$. Recalling the construction of the supersolution $u^{+}$ in Lemma 3.8, we see that $u^{+}\geq 0$ in $\Omega$. However, the subsolution $u^{-}$ defined in Lemma 3.8 is not necessarily nonnegative.
Now, assume $f=0$ on $\mathbb{R}^{-}$. Then, there will always be a nonnegative subsolution $u^{-}$ such that $u^{-}\leq u^{+}$ on $\overline{\Omega}$ and $u^{+}=u^{-}=g$ on $\partial\Omega$. In fact, it is easy to see that $w^{\star}:=\max\{u^{-},0\}$ is also a subsolution with $w^{\star}=g$ on $\partial\Omega$. Moreover, we have  $w^{\star}\leq u^{+}$. Hence, $w^{\star}\in S$. From the definition of the Perron’s solution $u$ (see Proposition 3.10), this yields that
@@
Finally, assume that there is a point $x_{0}\in\Omega$ such that $u(x_{0})=0$. Since $u\neq 0$ and $u\in C(\overline{\Omega})$, then there is a point $x^{\star}\neq x_{0}\in\Omega$ such that $u>\frac{u(x^{\star})}{2}>0$ on $B(x^{\star},\varepsilon)$, where $\varepsilon>0$ is small enough. Now, let $\varphi\in C^{1}(\Omega)\cap C(\overline{\Omega})$ be such that $\mbox{supp}(\varphi)\subset B(x^{\star},\varepsilon)$ and $0\leq\varphi\leq\frac{u(x^{\star})}{4}$. In particular, we have $u\geq\varphi$ and $u(x_{0})=\varphi(x_{0})=0$. Therefore, we must have
@@
Notice also that when $f=0$ on $\mathbb{R}^{-}$, we obtain from Proposition 3.9 the following uniform (does not depend on the solution $u$) estimate:
$||u||_{C^{0,\beta}(\overline{\Omega})}\leq C\bigg{(}\alpha,\,\beta,\,\mbox{diam}(\Omega),\,||g||_{C^{0,\beta}(\partial\Omega)},\,f(||g||_{\infty})\bigg{)}.$
@@ 4. Obstacle problem
In this section, we assume that $f:[0,\infty)\mapsto\mathbb{R}$ is continuous, non-negative and non-decreasing and $g\geq 0$ on $\partial\Omega$. Assuming $\alpha<1$. We prove Theorem 1.2 by showing that there exists a non-negative function $u$ that is locally $\alpha-$Hölder continuous and in $C^{0,\beta}(\overline{\Omega})$ for any $\beta<\alpha$, solution to the following obstacle problem
$\begin{cases}L[u]=f(u)\qquad&\text{in\,\, $\{u>0\}$}\\ u=g\qquad&\text{on \,\,$\partial\Omega$}\end{cases}.$ (4.0.1)
@@
In the case when $f(0)=0$, we extend $f$ by 0 on $\mathbb{R}^{-}$. Then, this extension (we still denote it by $f$) is continuous and non-decreasing on $\mathbb{R}$ and so, by Proposition 3.10, Problem (1.0.8) has a solution $u$. Thanks to Remark 3.11, $u>0$. Hence, $u$ solves Problem (4.0.1).
We consider now the case when $f(0)>0$. Let $f_{\varepsilon}$ be a sequence of non-decreasing continuous functions such that $f_{\varepsilon}=0$ on $\mathbb{R}^{-}$ and $f_{\varepsilon}=f$ on $[\varepsilon,+\infty)$. For every $\varepsilon>0$, by Proposition 3.10, we know that there exists a solution $u_{\varepsilon}$ to Problem (1.0.8) with  $u_{\varepsilon}=g$  on  $\partial\Omega$. Recalling Remark 3.11, we may assume that $u_{\varepsilon}>0$ on $\Omega$. In addition, by Proposition 3.9, we have that
$||u_{\varepsilon}||_{C^{0,\beta}(\overline{\Omega})}\leq C\bigg{(}\alpha,\,\beta,\,\mbox{diam}(\Omega),\,||g||_{C^{0,\beta}(\partial\Omega)},\,f(||g||_{\infty})\bigg{)},$
for $\varepsilon>0$ small enough. Hence, $(u_{\varepsilon})_{\varepsilon}$ is bounded in $C^{0,\beta}(\overline{\Omega})$. Therefore, up to a subsequence, $u_{\varepsilon}\rightarrow u$ uniformly in  $C^{0,\beta}(\overline{\Omega})$  and  $u\geq 0$  on  $\overline{\Omega}$.
@@
$-L[\varphi_{\varepsilon}](x_{\varepsilon})+f_{\varepsilon}(\varphi_{\varepsilon}(x_{\varepsilon}))\leq 0.$ (4.0.2)
@@
$|L[\varphi_{\varepsilon}](x_{\varepsilon})-L[\varphi](x_{\varepsilon})|\leq C||\varphi_{\varepsilon}-\varphi||_{\infty}^{1-\alpha}.$
@@
$\dfrac{\varphi_{\varepsilon}(y)-\varphi(y)+\varphi(x_{\varepsilon})-\varphi_{\varepsilon}(x_{\varepsilon})}{|y-x_{\varepsilon}|^{\alpha}}\leq 2\,\min\bigg{\{}\dfrac{||\varphi_{\varepsilon}-\varphi||_{\infty}}{|y-x_{\varepsilon}|^{\alpha}},C\,|y-x_{\varepsilon}|^{1-\alpha}\bigg{\}},$
@@
$\dfrac{\varphi_{\varepsilon}(y)-\varphi(y)+\varphi(x_{\varepsilon})-\varphi_{\varepsilon}(x_{\varepsilon})}{|y-x_{\varepsilon}|^{\alpha}}\leq C\,\min\bigg{\{}\dfrac{||\varphi_{\varepsilon}-\varphi||_{\infty}}{|y-x_{\varepsilon}|^{\alpha}},\,|y-x_{\varepsilon}|^{1-\alpha}\bigg{\}}\leq C||\varphi_{\varepsilon}-\varphi||_{\infty}^{1-\alpha}.$
@@
Finally, following the same argument in Proposition 3.9 and using the fact that $f=0$ on $\mathbb{R}^{-}$ and Propositon 2.3, we get that $u_{\epsilon}$ are locally $\alpha-$Hölder continuous uniformly in $\epsilon$, then letting $\epsilon\to 0$ we get that $u$ is locally Hölder continuous. ∎

### v2

@@ 3.2. Perron’s Method
$\limsup_{n\to\infty}u^{-}(y_{n})\leq u^{-}(y_{0}).$
Assume now that $b_{n}$ has a bounded subsequence (we still call it $b_{n}$). Yet, we note that thanks to (3.2.3) and the fact that $u^{-}$ is bounded from below and $a_{n}\leq\max g$, then $a_{n}$ is always bounded. So, we can extract subsequences, $(x_{n},a_{n},b_{n})\to(x_{1},a_{1},b_{1})\in\partial\Omega\times\mathbb{R}\times[0,\infty).$ From (3.2.3),
$u^{-}(y_{n})\leq a_{n}-b_{n}|y_{n}-x_{n}|^{\beta}+\dfrac{1}{n}.$
So,
$\limsup_{n\to\infty}u^{-}(y_{n})\leq a_{1}-b_{1}|y_{0}-x_{1}|^{\beta}=\lim_{n\to\infty}a_{n}-b_{n}|y_{0}-x_{n}|^{\beta}=\lim_{n\to\infty}\phi_{x_{n},a_{n},b_{n}}(y_{0})\leq u^{-}(y_{0}).$
We deduce that $u^{-}\in C(\overline{\Omega})$ and so we can use the stability result in Proposition 3.6 to conclude that $u^{-}$ is a subsolution to (1.0.8) and $u^{-}=g$ on $\partial\Omega.$
Similarly, we define
$S^{\prime}=\bigg\{(x_{0},a,b)\in\partial\Omega\times\mathbb{R}\times[0,\infty):L[\phi_{x_{0},a,-b}]\leq f(\phi_{x_{0},a,-b})\,\,\,\,\text{in\,\, $\Omega$},\,\,{\phi_{x_{0},a,-b}}\geq g\,\,\,\mbox{on}\,\,\,\partial\Omega\bigg\}.$
Using the same approach, one can show that $u^{+}=\inf\limits_{(x_{0},a,b)\in S^{\prime}}\phi_{x_{0},a,-b}$ is in $C(\overline{\Omega})$ and is a supersolution to (1.0.8) with $u^{+}=g$ on $\partial\Omega$.
Finally, we show that $u^{-}\leq u^{+}.$ Take $(x_{0},a_{0},b_{0})\in S$ and $(x_{0}^{\prime},a_{0}^{\prime},b_{0}^{\prime})\in S^{\prime}$. Since $\beta<1$ then $\phi_{x_{0},a_{0},b_{0}}-\phi_{x_{0}^{\prime},a_{0}^{\prime},-b_{0}^{\prime}}$ is a convex function over the compact domain $\overline{\Omega}$, then its global maximum is attained on $\partial\Omega$. But by definition of $S$ and $S^{\prime}$, $\phi_{x_{0},a_{0},b_{0}}\leq g\leq\phi_{x_{0}^{\prime},a_{0}^{\prime},-b_{0}^{\prime}}$ on $\partial\Omega$ concluding that $\phi_{x_{0},a_{0},b_{0}}\leq\phi_{x_{0}^{\prime},a_{0}^{\prime},-b_{0}^{\prime}}$ on $\overline{\Omega}$. Yet, this is true for every $(x_{0},a_{0},b_{0})\in S$ and $(x_{0}^{\prime},a_{0}^{\prime},b_{0}^{\prime})\in S^{\prime}$. Hence, $u^{-}\leq u^{+}$.
Remark 3.9.
Assume $g\geq 0$ on $\partial\Omega$. Then, for $(x_{0},a_{0},b_{0})\in S^{\prime}$, we have $\phi_{x_{0},a_{0},-b_{0}}\geq a_{0}=\phi_{x_{0},a_{0},-b_{0}}(x_{0})\geq g(x_{0})\geq\min g$. Hence, we get that $u^{+}\geq 0$; this observation will be needed in Section 4.
Now, we are ready to prove Theorem 1.1. First, we show the regularity of subsolutions and then prove the existence result.
@@
Assume $f$ is non-decreasing and continuous. Let  $u$  be a bounded viscosity subsolution of (1.0.8). Then, $u$ is locally $\beta$-Hölderian, for any $0<\beta<\alpha$. More precisely, we have
[…]
@@
In addition, assume that $g\in C^{0,\beta}(\partial\Omega)$, if $u$ is a bounded viscosity solution of (1.0.8) with $u=g$ on $\partial\Omega$, then $u$ is $\beta$-Hölderian in $\overline{\Omega}$, and we have the following estimate:
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Proposition 3.11.
Assume  $0<\alpha<1$, $f\geq 0$ and $f$ is non-decreasing and continuous, then any bounded viscosity subsolution $u$ of (1.0.8) is locally $\alpha$-Hölderian. In addition, we have
[…]
Proof.
We will follow the same argument as in Proposition 3.10. Let $\omega\subset\subset\Omega$ and fix $x_{0}\in\omega$. Set
$v(x)=u(x_{0})+C|x-x_{0}|^{\alpha}.$
Thanks to Propositon 2.3 and the fact that $f\geq 0$, one has
$L[u(x_{0})+C|x-x_{0}|^{\alpha}]+f(u(x_{0})+C|x-x_{0}|^{\alpha})\geq-CL[|x-x_{0}|^{\alpha}]>0.$
[…]
$u(x)\leq u(x_{0})+C|x-x_{0}|^{\alpha}\qquad\mbox{for all}\,\,\,\,x\in\omega.\qed$
@@
Under the assumptions that $f$ is non-decreasing and continuous, and $g$ is continuous on $\partial\Omega$, there exists a viscosity solution $u$ to Problem (1.0.8).
@@
From Lemma 3.8, we know that there exist a subsolution $u^{-}$ and a supersolution $u^{+}$ to (1.0.8) with $^{\pm}\in C(\overline{\Omega}), uu$$^{-}\leq u^{+}$ on  $\overline{\Omega}$  and  $u^{+}=u^{-}=g$  on  $\partial\Omega$. Set
@@
[…]
On the other hand, by Proposition 3.10 , one has
$w]_{C^{0,\beta}(\omega)}\leq\max\bigg\{\frac{2\Lambda}{\text{dist}(\omega,\partial\Omega)^{\beta}},\frac{[\mbox{diam}(\Omega)]^{\alpha-\beta}\,[f(-\Lambda)]_{-}}{1-\Psi(r_{\star})}\bigg\},\qquad\mbox{for every}\,\,\,w\subset\subset\Omega.$
Hence, we infer that $u$ is locally $\beta$-Hölder as it is the supremum of uniformly locally $\beta$-Hölder functions. So in particular, $u$ is continuous in $\Omega$. Yet, $u$ is continuous on $\partial\Omega$ since $u^{-}\leq u\leq u^{+}$ on $\overline{\Omega}$, $u^{\pm}\in C(\overline{\Omega})$ and, $u^{+}=u^{-}=g$ on $\partial\Omega$. From Proposition 3.6, this implies that $u$ is a subsolution of (1.0.8) with $u=g$ on $\partial\Omega$.
@@
$L[\varphi_{\delta}](x_{0})+f(\varphi_{\delta}(x_{0}))\leq-L[\varphi](x_{0})+C\delta+f(\varphi(x_{0}))<0,$
@@
$L[\varphi_{\zeta}](x_{\zeta})+f(\varphi_{\zeta}(x_{\zeta}))>0.$ (3.2.5)
@@
$L[\varphi_{\zeta}](x_{\zeta})+f(\varphi_{\zeta}(x_{\zeta}))\leq 0,$
@@
$L[\varphi_{\zeta}](x_{\zeta})+f(\varphi(x_{\zeta}))\leq-L[\varphi](x_{\zeta})+f(\varphi(x_{\zeta})).$
@@
$L[\varphi_{\zeta}](x_{\zeta})+f(\varphi_{\zeta}(x_{\zeta}))]+[f(\varphi(x_{\zeta}))-f(\varphi_{\zeta}(x_{\zeta}))]\leq-L[\varphi](x_{\zeta})+f(\varphi(x_{\zeta})).$ (3.2.6)
@@
Assume the boundary datum $g\geq 0$ on $\partial\Omega$ (but, $g$ is not identically zero). Recalling Remark 3.9, the supersolution $u^{+}$ constructed in Lemma 3.8 is nonnegative. However, the subsolution $u^{-}$ defined in Lemma 3.8 is not necessarily nonnegative.
Now, assume $f=0$ on $(-\infty,0]$}}. Then, there will always be a nonnegative subsolution $u^{-}$ such that $u^{-}\leq u^{+}$ on $\overline{\Omega}$ and $u^{+}=u^{-}=g$ on $\partial\Omega$. In fact, it is easy to see that $w^{\star}:=\max\{u^{-},0\}$ is also a subsolution with $w^{\star}=g$ on $\partial\Omega$. Moreover, we have  $w^{\star}\leq u^{+}$. Hence, using the setting of Proposition 3.12, $w^{\star}\in S$. From the definition of the Perron’s solution $u$, this yields that
@@
Finally, assume that there is a point $x_{0}\in\Omega$ such that $u(x_{0})=0$. Since $u\neq 0$ and $u\in C(\overline{\Omega})$, then there is a point $x^{\star}\neq x_{0}\in\Omega$ such that $u>\frac{u(x^{\star})}{2}>0$ on $B(x^{\star},\varepsilon)$, where $\varepsilon>0$ is small enough. Now, let $\varphi\in C^{1}(\Omega)\cap C(\overline{\Omega})$ be such that $\varphi\neq 0$, $\mbox{supp}(\varphi)\subset B(x^{\star},\varepsilon)$ and $0\leq\varphi\leq\frac{u(x^{\star})}{4}$. In particular, we have $u\geq\varphi$ and $u(x_{0})=\varphi(x_{0})=0$. Therefore, we must have
@@
Notice also that when $f=0$ on $(-\infty,0]$}}, we obtain from Proposition 3.10 the following uniform (does not depend on the solution $u$) estimate:
[…]
@@ 4. Obstacle problem
In this section, we assume that $f:[0,\infty)\mapsto\mathbb{R}$ is continuous, nonnegative and non-decreasing, and the boundary datum $g$ is nonnegative and continuous on $\partial\Omega$. We prove Theorem 1.2 by showing that there exists a nonnegative function $u$ that is solution to the following obstacle problem
$\begin{cases}L[u]=f(u)\qquad&\text{in\,\, $\{u>0\}$},\\ u=g\qquad&\text{on \,\,$\partial\Omega$}.\end{cases}$ (4.0.1)
@@
In the case when $f(0)=0$, we extend $f$ by 0 on $(-\infty,0)$}}. Then, this extension (we still denote it by $f$) is continuous and non-decreasing on $\mathbb{R}$ and so, by Proposition 3.12, Problem (1.0.8) has a solution $u$. Thanks to Remark 3.13, $u>0$. Hence, $u$ solves Problem (4.0.1).
Now, we consider the case when $f(0)>0$. Assuming $\alpha<1$. Let $f_{\varepsilon}$ be a sequence of non-decreasing continuous functions such that $f_{\varepsilon}=0$ on $(-\infty,0]$}} and $f_{\varepsilon}=f$ on $\varepsilon,+\infty)$. For every $\varepsilon>0$, by Proposition 3.12, we know that there exists a solution $u_{\varepsilon}$ to Problem (1.0.8) with  $u_{\varepsilon}=g$  on  $\partial\Omega$. Recalling Remark 3.13, we may assume that $u_{\varepsilon}>0$ on $\Omega$. In addition, by Proposition 3.10, we have that
[…]
for $\varepsilon>0$ small enough.
Hence, $(u_{\varepsilon})_{\varepsilon}$ is bounded in $C^{0,\beta}(\overline{\Omega})$. Therefore, up to a subsequence, $u_{\varepsilon}\rightarrow u$ uniformly in  $C^{0,\beta}(\overline{\Omega})$  and  $u\geq 0$  on  $\overline{\Omega}$.
@@
$L[\varphi_{\varepsilon}](x_{\varepsilon})+f_{\varepsilon}(\varphi_{\varepsilon}(x_{\varepsilon}))\leq 0.$ (4.0.2)
@@
[…]
@@
[…]
@@
[…]
@@
Thanks to Proposition 3.11 and since $f_{\varepsilon}\geq 0$, then $u_{\epsilon}$ are uniformly (in $\epsilon$) locally $\alpha$-Hölder continuous. Moreover, one has
[…]
Then, letting $\epsilon\to 0$, we infer that the limit function $u$ is locally $\alpha$-Hölder continuous. ∎
Remark 4.1.
We note that it is not clear whether we can prove global $\alpha$-Hölder regularity on the solution $u$ in the previous theorem or not, unless $f=0$. Indeed, recalling the proof of Proposition 3.10, the key idea is to show that, at each boundary point $x_{0}\in\partial\Omega$, the function $u_{\varepsilon}$ admits two barrier functions from above and below of the form $w^{\pm}(x)=g(x_{0})\pm C\,|x-x_{0}|^{\alpha}$. The existence of a barrier function $w^{+}$ from above is not an issue (see the proof of Proposition 3.10). However, the difficulty arises when showing a barrier function $w^{-}$ from below since now (i.e. when $\beta=\alpha$) we lose the upper bound $L[|x-x_{0}|^{\beta}]\leq-c<0$ (which is true if $\beta<\alpha$; see Proposition 2.2). In fact, set
$w^{-}(x)=g(x_{0})-C\,|x-x_{0}|^{\alpha}.$
We have
$L[w^{-}]+f(w^{-})=C\,L[|x-x_{0}|^{\alpha}]+f(g(x_{0})-C\,|x-x_{0}|^{\alpha}).$
By Proposition 2.3, we get
$L[w^{-}]+f(w^{-})\leq-1+\dfrac{\left(\dfrac{\text{{\rm diam}}(\Omega)}{|x-x_{0}|}\right)^{\alpha}-1}{\left(\dfrac{\text{{\rm diam}}(\Omega)}{|x-x_{0}|}-1\right)^{\alpha}}+f(g(x_{0})-C\,|x-x_{0}|^{\alpha}).$
But, the right hand side in the inequality above is not strictly negative as soon as $x$ is close enough to $x_{0}$ and $f(g(x_{0}))>0$. Hence, it is not clear whether $w^{-}$ is a strict subsolution to (1.0.8) or not.
Remark 4.2.
We note that the solution $u$ of the obstacle problem (4.0.1) is not necessarily strictly positive. Thus, an interesting question will be to study the properties of the free boundary $\partial\{u>0\}$, this remains open.
On the other hand, regularity results of the solution on the free boundary for the same problem (4.0.1) but with infinity (instead of fractional) Laplacian has been analyzed for example in [13] through a scaling argument that could not be adapted to our case since the operator $L$ is non-local. Finally, we mention that for the infinity Laplacian we obtain higher regularity on the solution across the free boundary. However, it is not clear if this will be also the case when dealing with fractional infinity Laplacian.
Acknowledgement
The authors would like to thank the referee for all their comments and in particular for their suggestions to refine our results under weaker regularity assumptions on the boundary datum $g$.